OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8

OpenStudy (anonymous):

Ax+By=C

OpenStudy (anonymous):

that is the standard form, the slope int form is y=mx+b

OpenStudy (anonymous):

None of these are right because I have A.-D. choices, and none say these.

OpenStudy (anonymous):

Slope-intercept form is y = mx + b. You have to simplify the equation you've been given to fit this.

OpenStudy (anonymous):

well you have to chance your equation to look like the slope-intercept equation: y=mx+b. so first isolate the y, then divide both sides by 4, that will be your answer.

OpenStudy (anonymous):

I divide the 6 and the 8 by 4? Or 6, 8, and 4.

OpenStudy (anonymous):

You need to get y by itself. First you need to move x to the other side, then you'd divide by what is attached to y so it's alone.

OpenStudy (anonymous):

so first bring everything to one side except the y. 4y=8-6x or 4y=-6x+8. then divide the 6 and 8 by 4 to make y all alone.

OpenStudy (anonymous):

remember to simplify the fractions! then that will be your answer